OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (anonymous):

Have you tried using point slope form? y2-y1=m(x2-x1) Once you have the equation in this form, you can use algebra to manipulate it into slope-intercept form (y=mx+b). Let me know if you need anything else! :)